OpenStudy (anonymous):

What is the Celsius temperature of 1 mole of a gas that has an average kinetic energy of 4,053 joules?

OpenStudy (cuanchi):

do you have any approach to solve the problem? you can look here http://hyperphysics.phy-astr.gsu.edu/hbase/kinetic/kintem.html#c1

OpenStudy (cuanchi):

for one mole of gas you can use the following equation, isolate T and then convert T from Kelvin to Celsius. Kinetic energy \[KE =\frac{ 3 }{ 2 } \times R \times T\] were R= 8.314 J/mol K and T= temperature in Kelvin